# Diffusion Models in Simulation-Based Inference: A Tutorial Review

### v1

@@ 5.3 Weighting Function
The weighting function $\omega(t)$ in score-matching (10) controls how much emphasis different noise levels receive during training. By scaling the contribution of each time step $t$ and thereby the corresponding $\lambda_{t}$, it biases optimization toward regions of the signal-to-noise ratio where improvements are expected to matter most (Figure 8). While some weighting rules have theoretical motivation (for example, ensuring consistency with variational bounds [SongDur2021]), most are empirical design choices tuned for better convergence, stability, or perceptual quality. Here, we present the weighting functions for the noise-prediction loss (58). A widely used choice is likelihood weighting [SongDur2021], which sets
@@
In contrast, EDM weighting [karras2022elucidating] uses
@@
Another option is sigmoid weighting [KingmaGao2023],
@@
In practice, most SBI papers either omit explicit weighting, effectively setting $\omega(t)=1$, or apply EDM weighting in combination with the EDM schedule (LABEL:tab:papers). In our experiments (Section 6), we find the EDM weighting to be superior with regard to the statistical accuracy of the diffusion estimators.
@@ 5.4 Model Parameterization
Diffusion models can be parameterized in terms of what they predict, and each choice can be advantageous at different noise levels. As we will see next, we can easily convert one parameterization into another. In score prediction [SongErm2020a], the network predicts the score $\hat{\boldsymbol{s}}$ directly. This parameterization is statistically elegant because the score function uniquely characterizes the target distribution. It has also been the most common choice in SBI, often used without explicit reweighting, which implicitly puts more weight on time steps in the low-noise regime. However, the network has to learn scores whose magnitudes vary by several orders across noise levels, which can make training difficult. To mitigate this, [song2020improved] proposed using a rescaled version of the score, $\boldsymbol{s}/\sigma_{t}$, as the training target, which normalizes the scale of the target across time steps and leads to more stable training across the full range of noise levels.
In noise prediction [HoJai2020a], the network instead predicts
@@
An alternative is target prediction [song2020denoising], where the diffusion model predicts the clean target
@@
In velocity prediction [salimans2022progressive], the model predicts a linear combination of $\boldsymbol{\epsilon}_{t}$ and $\mathbf{z}_{0}$:
@@
Finally, EDM prediction [karras2022elucidating] generalizes this idea by reconditioning the target with
@@
Each parameterization shifts the learning emphasis: $\boldsymbol{\epsilon}$-prediction is robust but weak at very low noise; $\mathbf{z}_{0}$-prediction excels at low noise but becomes unstable at high noise; velocity and EDM prediction recondition the problem to achieve uniform difficulty across time steps. While most SBI papers still use the score prediction, there is growing interest in EDM-style targets due to their improved stability and efficiency (LABEL:tab:papers). Crucially, with appropriate weighting of the loss, the different parameterizations are mathematically equivalent [KingmaGao2023]:
@@
If no correction is applied, switching the parameterization implicitly changes the weighting function. In fact, score prediction without weighting corresponds to using noise prediction with weights proportional to $1/\sigma_{t}^{2}$. A practical strategy is to predict any convenient target, map it back to the noise domain, and compute the loss there [salimans2022progressive]. This ensures that the parameterization does not distort the effective weighting over noise levels $\lambda_{t}$. Different methods, such as continuous normalizing flows, flow matching, and consistency models, are all connected to diffusion models by the choice of parameterization of the neural estimator, as will be discussed in Section 5.6.
@@ 5.5 Design Choices During Inference
For a known forward process, the drift $f(t)$ and the diffusion coefficient $g(t)$ can be expressed through the log signal-to-noise ratio (SNR) $\lambda_{t}$ [KingmaGao2023]. In the variance-preserving case, the drift and the diffusion coefficient are given by
@@
with $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(\boldsymbol{0},\mathbf{I})$. This continuous-time scheme recovers the discrete sampler of [HoJai2020a].
This stochastic formulation permits the use of a broad range of SDE solvers. Because the dynamics include an additive noise term, one can employ solvers that achieve higher accuracy than the standard Euler–Maruyama scheme. Examples include SEA, a one-step method with strong order 1 for additive-noise SDEs [foster2024high], and ShARK, a two-step method attaining strong order 1.5 [foster2024high]. In our experiments (Section 6.1), the adaptive two-step solver introduced by [Jolicoeur-MartineauLi2021] provides the most favorable balance between speed and accuracy.
@@
We have seen that every diffusion SDE has an associated deterministic process called the probability flow ODE [SongSoh2020]. The key observation is that the Fokker–Planck equation governing the marginal densities of the SDE admits both a stochastic solution (the reverse-time SDE) and a deterministic solution (the ODE). Both processes share the same marginal distributions $p(\mathbf{z}_{t})$ for all $t\in[0,1]$, meaning that solving either produces identically distributed samples in theory. The probability flow ODE can be expressed as
@@
The ODE provides a direct flow that transports the distribution smoothly from randomly sampled noise at $t=1$ to the target at $t=0$. In terms of $\alpha_{t}$ and $\sigma_{t}$, [karras2022elucidating] express the ODE as
@@
However, ODE solvers can potentially suffer from collapse errors, since deterministic flows may fail to explore the typical set of $p(\mathbf{z}\mid\mathbf{x})$. In contrast, the stochasticity in SDE solvers naturally spreads probability mass more broadly. Recent work shows that network architectures with skip connections (e.g., EDM parameterization) help mitigate mode collapse [zhang2024collapse]. Our experiments suggest that SDE solvers outperform ODE solvers slightly on low-dimensional problems, but not on high-dimensional ones (Section 6). In SBI, collapse errors can be diagnosed by checking the calibration of the approximate posteriors [modrak2025simulation].
@@
Besides solving either the ODE or SDE, one can also employ the learned score directly in annealed Langevin dynamics [SongErm2020a]. Langevin sampling evolves target samples through a discretized stochastic gradient ascent on the log density,
@@
In contrast to solving the reverse SDE or ODE, the Langevin formulation explicitly discretizes time and applies stochastic updates at each step. This can be useful if the score is modified post hoc after training, since the corresponding marginal distribution $p(\mathbf{z}_{t}\mid\mathbf{z}_{0})$ may then belong to a different reverse process [geffner2023compositional, linhart2024diffusion, arruda2025compositional]. However, in practice, the reverse SDE is often still employed using a modified score with good results (Section 4). [SongSoh2020] introduce corrector–predictor sampling, which combines an Euler–Maruyama step with an additional annealed Langevin step to improve image quality. However, our results do not show an increased statistical accuracy for SBI tasks (Section A.1.1).
So far, most SBI papers have employed either continuous or discretized versions of the reverse ODE or SDE for inference (LABEL:tab:papers). In the SBI setting, posterior calibration can be systematically checked using simulation-based calibration [talts2018validating, modrak2025simulation, SBC;], making it straightforward to compare samplers in terms of statistical accuracy and efficiency. Since the choice of sampler is made at inference time, it does not affect training. This flexibility allows users to switch between different solvers depending on whether the goal is fast point estimation, fully Bayesian inference, or likelihood evaluation.
@@ 5.6 Design Choices Beyond the SDE Paradigm
Diffusion models are most commonly expressed through the score and the corresponding SDE, where noisy trajectories are simulated forward and reversed during inference (Section 3). However, recent work has introduced alternative training objectives that bypass explicit SDE simulation. These methods aim to retain the flexibility of diffusion models, but simplify parameterization and accelerate inference. Two prominent model families have emerged as particularly relevant for SBI: conditional flow matching [wildberger2023flow] and conditional consistency models [schmitt2024consistency]. Both can be interpreted as competitive extensions or simplifications of score matching that do not require solving an SDE for sampling.
@@ 5.6.1 Conditional Flow Matching
This formulation corresponds to the conditional version of flow matching [lipman2023flow], also introduced as rectified flow [liu2023flow] and discussed by [lu2025simplifying].
@@
[KingmaGao2023] show that the loss defined on the vector field $v(\mathbf{z}_{t},\mathbf{x},t)$ is equivalent to the standard noise-prediction loss, up to a time-dependent weighting:
@@
In practice, flow matching is usually trained with the unweighted objective on $v$ and with the simple schedule $\alpha_{t}=1-t$, $\sigma_{t}=t$, where $t\sim\mathcal{U}[0,1]$. Flow matching tends to produce straighter trajectories compared to standard score-based parameterizations, and hence allows faster simulation of the corresponding ODE [liu2023flow].
In the context of SBI, conditional flow matching was applied by [wildberger2023flow], who also proposed a power law distribution for $t$, with $p(t)\propto(1-t)^{1/(1+\rho)}$ and $\rho>-1$. Using this distribution corresponds to changing the weighting function, that is, assigning greater importance to the vector field closer to $\mathbf{z}_{0}$ for $\rho<0$. [wildberger2023flow] and [orsini2025flow] empirically found that $\rho<0$ improves learning for distributions with sharp bounds. We observed that a value of $\rho=-0.6$ gives a weighting distribution of $\lambda_{t}$ similar to the EDM formulation (Figure 7D) and that it improves learning only on low-dimensional SBI tasks (Section 6). Similarly, [gebhard2025flow] put more emphasis on higher signal-to-noise ratio by sampling $\sqrt[4]{t}\sim\mathcal{U}[0,1]$.
@@
An important extension is optimal transport conditional flow matching (OT-CFM), proposed by [tong2024improving] and [pooladian2023multisample]. This approach generalizes standard flow matching by replacing the independent coupling $(\mathbf{z}_{0},\mathbf{z}_{1})$ with the 2-Wasserstein optimal transport plan $\pi(\mathbf{z}_{0},\mathbf{z}_{1})$ [benamou2000computational]. Instead of sampling $\mathbf{z}_{0}$ and $\mathbf{z}_{1}$ independently, one samples $(\mathbf{z}_{0},\mathbf{z}_{1})\sim\pi$, which encourages interpolation paths to respect the optimal transport geometry of the target distribution. This modification has a key advantage: learned flows are straighter (Figure 7E), since trajectories follow optimal transport geodesics. Hence, inference is accelerated, as the network learns paths that are naturally shorter and closer to the target manifold. This seems to be particularly helpful for high-dimensional inference problems (Section 6).
The main drawback of OT-CFM is computational: training typically takes about $2.5\times$ longer due to the cost of computing and sampling from optimal transport couplings. The latter requires large batch sizes to find better pairs [zhang2025fittingflowmodelslarge], while learned couplings might be optimal on the batch level but partially wrong in comparison with the optimal transportation plan [nguyen2022improving]. Recent work tries to overcome the computational bottleneck with a semi-discrete formulation [mousavi2025flow]. Moreover, [Fluri2024-improvingFlowMatchingSimulation-basedInferenceu] and [cheng2025curse] adapted this idea to the conditional setting. Since standard mini-batch reordering ignores the condition $\mathbf{x}$, the authors propose a modified transport cost with an additional penalty term $d(\mathbf{x}^{(i)},\mathbf{x}^{(j)})$ leading to conditional optimal transport (COT), where $d$ is a weighted distance measure on the raw or embedded observations $\mathbf{x}$ (Section A.1.1).
@@ 5.6.2 Conditional Consistency Models
Consistency models in discrete and continuous time [song2023consistency] can be seen as a distillation scheme for diffusion models, or as a standalone generative model family. In the context of SBI, discrete-time conditional consistency models were adapted by [schmitt2024consistency]. These models learn a consistency function $c$ that maps a noisy state $\mathbf{z}_{t}$ directly to the clean target $\mathbf{z}_{0}$. The defining property is that the mapping should be consistent across all points $(\mathbf{z}_{t},t)$ along the same probability flow ODE trajectory: $c(\mathbf{z}_{t},\mathbf{x},t)=c(\mathbf{z}_{t^{\prime}},\mathbf{x},t^{\prime})$, with boundary condition $c(\mathbf{z}_{0},\mathbf{x},0)=\mathbf{z}_{0}$. In practice, the consistency function is parameterized using skip connections:
@@
There are two principal approaches to training consistency models. The first distills a pre-trained diffusion model by sampling points along its probability flow ODE trajectory (8) and using them as supervision. This method is computationally expensive, as it requires repeated sampling from the diffusion model, and has not been applied to SBI. The second approach, adopted by [schmitt2024consistency], trains a consistency model directly as a standalone estimator without relying on a pre-trained diffusion model. The training objective for discrete conditional consistency models is
@@
where $\Delta t>0$ is the time step size, $\omega_{t}$ is a weighting function, $d$ is a distance metric, and $\bar{c}(\cdot)$ denotes a teacher model that is an exponential moving average (EMA) version of the student network $c(\cdot)$. [song2023improved] suggested to use $\omega_{t}=1/\Delta t$, a Pseudo-Huber distance, a curriculum for the discretization steps, and to update the student model directly with the parameters of the teacher when trained in isolation. These settings were adopted by [schmitt2024consistency] for SBI, together with a variance-exploding EDM noise schedule.
@@
which is sufficient when using Euler integration as the ODE solver [song2023consistency]. This substitution allows consistency models to be trained without explicit reliance on diffusion models, making them a flexible standalone sampler for SBI. However, using an ODE solver introduces additional discretization error. To avoid this, [song2023consistency] shows that, in the limit $\Delta t\to 0$ with a squared error distance $d(\mathbf{z}_{0},\hat{\mathbf{z}}_{0})=\|\mathbf{z}_{0}-\hat{\mathbf{z}}_{0}\|_{2}^{2}$, the gradient of (73) can be computed directly, without explicitly solving the ODE. Building on this idea, [lu2025simplifying] propose continuous consistency models based on trigonometric noise schedules derived from the unit-variance principle (similar to the EDM schedule). Their consistency function uses $c_{\text{skip}}(t)=\cos(0.5\pi t)$ and $c_{\text{out}}(t)=-\sin(0.5\pi t)\sigma_{\text{target}}$ with target standard deviation $\sigma_{\text{target}}$, yielding an MSE objective whose gradient matches (73). To further stabilize training, [lu2025simplifying] introduce several improvements: tangent normalization, tangent warm-up, and adaptive weighting, for which we refer to the original publication.
@@
Repeating these steps for decreasing $\tau_{k}$ produces a refined trajectory of reconstructions, with the final output $\hat{\mathbf{z}}_{0}^{(\tau_{K})}$ serving as a sample from $q(\mathbf{z}\mid\mathbf{x})\approx p(\mathbf{z}\mid\mathbf{x})$. [schmitt2024consistency] show that few-step sampling in the order of $\approx 10$ steps strikes a good balance between accuracy and sampling speed on SBI problems.
In the SBI setting, consistency models have important tradeoffs. A key limitation is that they do not allow density evaluation at arbitrary targets $\mathbf{z}$, since the inverse of the consistency mapping is not available in closed form. However, they offer few-step sampling and strong performance in low-data regimes [schmitt2024consistency]. To the best of our knowledge, continuous consistency models [lu2025simplifying] had not been applied in SBI prior to this work. Discrete variants have appeared only in the original work and in a subsequent medical imaging study [zhao2025generativeconsistencymodelsestimation], where they matched the accuracy of diffusion models and MCMC, but achieved superior sampling speed.

### v2

@@ 5.3 Weighting Function
The weighting function $\omega(t)$ in score-matching (10) controls how much emphasis different noise levels receive during training. By scaling the contribution of each time step $t$ and thereby the corresponding $\lambda_{t}$, it biases optimization toward regions of the signal-to-noise ratio where improvements are expected to matter most (Figure 8). While some weighting rules have theoretical motivation (for example, ensuring consistency with variational bounds [163]), most are empirical design choices tuned for better convergence, stability, or perceptual quality. Here, we present the weighting functions for the noise-prediction loss (58). A widely used choice is likelihood weighting [163], which sets
@@
In contrast, EDM weighting [88] uses
@@
Another option is sigmoid weighting [92],
@@
In practice, most SBI papers either omit explicit weighting, effectively setting $\omega(t)=1$, or apply EDM weighting in combination with the EDM schedule (LABEL:tab:papers). In our experiments (Section 6), we find the EDM weighting to be superior for low-dimensional problems with regard to the statistical accuracy of the diffusion estimators, while for high-dimensional problems sigmoid weighting seems to be a bit more accurate.
@@ 5.4 Model Parameterization
Diffusion models can be parameterized in terms of what they predict, and each choice can be advantageous at different noise levels. As we will see next, we can easily convert one parameterization into another. In score prediction [164], the network predicts the score $\hat{\boldsymbol{s}}$ directly. This parameterization is statistically elegant because the score function uniquely characterizes the target distribution. It has also been the most common choice in SBI, often used without explicit reweighting, which implicitly puts more weight on time steps in the low-noise regime. However, the network has to learn scores whose magnitudes vary by several orders across noise levels, which can make training difficult. To mitigate this, [165] proposed using a rescaled version of the score, $\boldsymbol{s}/\sigma_{t}$, as the training target, which normalizes the scale of the target across time steps and leads to more stable training across the full range of noise levels.
In noise prediction [71], the network instead predicts
@@
An alternative is target prediction [160], where the diffusion model predicts the clean target
@@
In velocity prediction [147], the model predicts a linear combination of $\boldsymbol{\epsilon}_{t}$ and $\mathbf{z}_{0}$:
@@
Finally, EDM prediction [88] generalizes this idea by reconditioning the target with
@@
Each parameterization shifts the learning emphasis: $\boldsymbol{\epsilon}$-prediction is robust but weak at very low noise; $\mathbf{z}_{0}$-prediction excels at low noise but becomes unstable at high noise; velocity and EDM prediction recondition the problem to achieve uniform difficulty across time steps. While most SBI papers still use the score prediction, there is growing interest in EDM-style targets due to their improved stability and efficiency (LABEL:tab:papers). Crucially, with appropriate weighting of the loss, the different parameterizations are mathematically equivalent [92]:
@@
If no correction is applied, switching the parameterization implicitly changes the weighting function. In fact, score prediction without weighting corresponds to using noise prediction with weights proportional to $1/\sigma_{t}^{2}$. A practical strategy is to predict any convenient target, map it back to the noise domain, and compute the loss there [147]. This ensures that the parameterization does not distort the effective weighting over noise levels $\lambda_{t}$. Different methods, such as continuous normalizing flows, flow matching, and consistency models, are all connected to diffusion models by the choice of parameterization of the neural estimator, as will be discussed in Section 5.6.
@@ 5.5 Design Choices During Inference
For a known forward process, the drift $f(t)$ and the diffusion coefficient $g(t)$ can be expressed through the log signal-to-noise ratio (SNR) $\lambda_{t}$ [92]. In the variance-preserving case, the drift and the diffusion coefficient are given by
@@
with $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(\boldsymbol{0},\mathbf{I})$. This continuous-time scheme recovers the discrete sampler of [71].
This stochastic formulation permits the use of a broad range of SDE solvers. Because the dynamics include an additive noise term, one can employ solvers that achieve higher accuracy than the standard Euler–Maruyama scheme. Examples include SEA, a one-step method with strong order 1 for additive-noise SDEs [48], and ShARK, a two-step method attaining strong order 1.5 [48]. In our experiments (Section 6.1), the adaptive two-step solver introduced by [87] provides the most favorable balance between speed and accuracy.
@@
We have seen that every diffusion SDE has an associated deterministic process called the probability flow ODE [166]. The key observation is that the Fokker–Planck equation governing the marginal densities of the SDE admits both a stochastic solution (the reverse-time SDE) and a deterministic solution (the ODE). Both processes share the same marginal distributions $p(\mathbf{z}_{t})$ for all $t\in[0,1]$, meaning that solving either produces identically distributed samples in theory. The probability flow ODE can be expressed as
@@
The ODE provides a direct flow that transports the distribution smoothly from randomly sampled noise at $t=1$ to the target at $t=0$. In terms of $\alpha_{t}$ and $\sigma_{t}$, [88] express the ODE as
@@
However, ODE solvers can potentially suffer from collapse errors, since deterministic flows may fail to explore the typical set of $p(\mathbf{z}\mid\mathbf{x})$. In contrast, the stochasticity in SDE solvers naturally spreads probability mass more broadly. Recent work shows that network architectures with skip connections (e.g., EDM parameterization) help mitigate mode collapse [203]. Our experiments suggest that SDE solvers outperform ODE solvers slightly on low-dimensional problems, but not on high-dimensional ones (Section 6). In SBI, collapse errors can be diagnosed by checking the calibration of the approximate posteriors [118].
@@
Besides solving either the ODE or SDE, one can also employ the learned score directly in annealed Langevin dynamics [164]. Langevin sampling evolves target samples through a discretized stochastic gradient ascent on the log density,
@@
In contrast to solving the reverse SDE or ODE, the Langevin formulation explicitly discretizes time and applies stochastic updates at each step. This can be useful if the score is modified post hoc after training, since the corresponding marginal distribution $p(\mathbf{z}_{t}\mid\mathbf{z}_{0})$ may then belong to a different reverse process [53, 105, 4]. However, in practice, the reverse SDE is often still employed using a modified score with good results (Section 4). [166] introduce corrector–predictor sampling, which combines an Euler–Maruyama step with an additional annealed Langevin step to improve image quality. However, our results do not show an increased statistical accuracy for SBI tasks (Section A.1.1).
So far, most SBI papers have employed either continuous or discretized versions of the reverse ODE or SDE for inference (LABEL:tab:papers). In the SBI setting, posterior calibration can be systematically checked using simulation-based calibration [169, 118, SBC;], making it straightforward to compare samplers in terms of statistical accuracy and efficiency. Since the choice of sampler is made at inference time, it does not affect training. This flexibility allows users to switch between different solvers depending on whether the goal is fast point estimation, fully Bayesian inference, or likelihood evaluation.
@@ 5.6 Design Choices Beyond the SDE Paradigm
Diffusion models are most commonly expressed through the score and the corresponding SDE, where noisy trajectories are simulated forward and reversed during inference (Section 3). However, recent work has introduced alternative training objectives that bypass explicit SDE simulation. These methods aim to retain the flexibility of diffusion models, but simplify parameterization and accelerate inference. Two prominent model families have emerged as particularly relevant for SBI: conditional flow matching [190] and conditional consistency models [151]. Both can be interpreted as competitive extensions or simplifications of score matching that do not require solving an SDE for sampling.
@@ 5.6.1 Conditional Flow Matching
This formulation corresponds to the conditional version of flow matching [106], also introduced as rectified flow [107] and discussed by [113].
@@
[92] show that the loss defined on the vector field $v(\mathbf{z}_{t},\mathbf{x},t)$ is equivalent to the standard noise-prediction loss, up to a time-dependent weighting:
@@
In practice, flow matching is usually trained with the unweighted objective on $v$ and with the simple schedule $\alpha_{t}=1-t$, $\sigma_{t}=t$, where $t\sim\mathcal{U}[0,1]$. Flow matching tends to produce straighter trajectories compared to standard score-based parameterizations, and hence allows faster simulation of the corresponding ODE [107].
In the context of SBI, conditional flow matching was applied by [190], who also proposed a power law distribution for $t$, with $p(t)\propto(1-t)^{1/(1+\rho)}$ and $\rho>-1$. Using this distribution corresponds to changing the weighting function, that is, assigning greater importance to the vector field closer to $\mathbf{z}_{0}$ for $\rho<0$. [190] and [128] empirically found that $\rho<0$ improves learning for distributions with sharp bounds. We observed that a value of $\rho=-0.6$ gives a weighting distribution of $\lambda_{t}$ similar to the EDM formulation (Figure 7D) and that it improves learning only on low-dimensional SBI tasks (Section 6). Similarly, [52] put more emphasis on higher signal-to-noise ratio by sampling $\sqrt[4]{t}\sim\mathcal{U}[0,1]$.
@@
An important extension is optimal transport conditional flow matching (OT-CFM), proposed by [175] and [137]. This approach generalizes standard flow matching by replacing the independent coupling $(\mathbf{z}_{0},\mathbf{z}_{1})$ with the 2-Wasserstein optimal transport plan $\pi(\mathbf{z}_{0},\mathbf{z}_{1})$ [15]. Instead of sampling $\mathbf{z}_{0}$ and $\mathbf{z}_{1}$ independently, one samples $(\mathbf{z}_{0},\mathbf{z}_{1})\sim\pi$, which encourages interpolation paths to respect the optimal transport geometry of the target distribution. This modification has a key advantage: learned flows are straighter (Figure 7E), since trajectories follow optimal transport geodesics. Hence, inference is accelerated, as the network learns paths that are naturally shorter and closer to the target manifold. This seems to be particularly helpful for high-dimensional inference problems (Section 6).
The main drawback of OT-CFM is computational: training typically takes about $2.5\times$ longer due to the cost of computing and sampling from optimal transport couplings. The latter requires large batch sizes to find better pairs [200], while learned couplings might be optimal on the batch level but partially wrong in comparison with the optimal transportation plan [122]. Recent work tries to overcome the computational bottleneck with a semi-discrete formulation [119]. Moreover, [47] and [26] adapted this idea to the conditional setting. Since standard mini-batch reordering ignores the condition $\mathbf{x}$, the authors propose a modified transport cost with an additional penalty term $d(\mathbf{x}^{(i)},\mathbf{x}^{(j)})$ leading to conditional optimal transport (COT), where $d$ is a weighted distance measure on the raw or embedded observations $\mathbf{x}$ (Section A.1.1).
@@ 5.6.2 Conditional Consistency Models
Consistency models in discrete and continuous time [161] can be seen as a distillation scheme for diffusion models, or as a standalone generative model family. In the context of SBI, discrete-time conditional consistency models were adapted by [151]. These models learn a consistency function $c$ that maps a noisy state $\mathbf{z}_{t}$ directly to the clean target $\mathbf{z}_{0}$. The defining property is that the mapping should be consistent across all points $(\mathbf{z}_{t},t)$ along the same probability flow ODE trajectory: $c(\mathbf{z}_{t},\mathbf{x},t)=c(\mathbf{z}_{t^{\prime}},\mathbf{x},t^{\prime})$, with boundary condition $c(\mathbf{z}_{0},\mathbf{x},0)=\mathbf{z}_{0}$. In practice, the consistency function is parameterized using skip connections:
@@
There are two principal approaches to training consistency models. The first distills a pre-trained diffusion model by sampling points along its probability flow ODE trajectory (8) and using them as supervision. This method is computationally expensive, as it requires repeated sampling from the diffusion model, and has not been applied to SBI. The second approach, adopted by [151], trains a consistency model directly as a standalone estimator without relying on a pre-trained diffusion model. The training objective for discrete conditional consistency models is
@@
where $\Delta t>0$ is the time step size, $\omega_{t}$ is a weighting function, $d$ is a distance metric, and $\bar{c}(\cdot)$ denotes a teacher model that is an exponential moving average (EMA) version of the student network $c(\cdot)$. [162] suggested to use $\omega_{t}=1/\Delta t$, a Pseudo-Huber distance, a curriculum for the discretization steps, and to update the student model directly with the parameters of the teacher when trained in isolation. These settings were adopted by [151] for SBI, together with a variance-exploding EDM noise schedule.
@@
which is sufficient when using Euler integration as the ODE solver [161]. This substitution allows consistency models to be trained without explicit reliance on diffusion models, making them a flexible standalone sampler for SBI. However, using an ODE solver introduces additional discretization error. To avoid this, [161] shows that, in the limit $\Delta t\to 0$ with a squared error distance $d(\mathbf{z}_{0},\hat{\mathbf{z}}_{0})=\|\mathbf{z}_{0}-\hat{\mathbf{z}}_{0}\|_{2}^{2}$, the gradient of (73) can be computed directly, without explicitly solving the ODE. Building on this idea, [113] propose continuous consistency models based on trigonometric noise schedules derived from the unit-variance principle (similar to the EDM schedule). Their consistency function uses $c_{\text{skip}}(t)=\cos(0.5\pi t)$ and $c_{\text{out}}(t)=-\sin(0.5\pi t)\sigma_{\text{target}}$ with target standard deviation $\sigma_{\text{target}}$, yielding an MSE objective whose gradient matches (73). To further stabilize training, [113] introduce several improvements: tangent normalization, tangent warm-up, and adaptive weighting, for which we refer to the original publication.
@@
Repeating these steps for decreasing $\tau_{k}$ produces a refined trajectory of reconstructions, with the final output $\hat{\mathbf{z}}_{0}^{(\tau_{K})}$ serving as a sample from $q(\mathbf{z}\mid\mathbf{x})\approx p(\mathbf{z}\mid\mathbf{x})$. [151] show that few-step sampling in the order of $\approx 10$ steps strikes a good balance between accuracy and sampling speed on SBI problems.
In the SBI setting, consistency models have important tradeoffs. A key limitation is that they do not allow density evaluation at arbitrary targets $\mathbf{z}$, since the inverse of the consistency mapping is not available in closed form. However, they offer few-step sampling and strong performance in low-data regimes [151]. To the best of our knowledge, continuous consistency models [113] had not been applied in SBI prior to this work. Discrete variants have appeared only in the original work and in a subsequent medical imaging study [204], where they matched the accuracy of diffusion models and MCMC, but achieved superior sampling speed.